The air in a hot-air balloon at 775 torr is heated from 23.0°C to 25.0°C. Assuming that the moles of air and the pressure remains constant, what is the density of the air at each temperature? (The average molar mass of air is 29.0 g/mol)

Density at 23.0°C
_____ g/L

density at 25.0°C
_____ g/L
Answer #1

n= constanta P= 775 torn = 1:02am (1 Torno 1.316X10-3 almy TL = 23°C: 273723 )<= 296k t2 = 25c =(273 +25) K = 2986 Main = 29